OpenStudy (anonymous):

if (x+5)(x+p)=x^2+2x+k, then ?

OpenStudy (phi):

first: use FOIL to multiply out the stuff on the left side

OpenStudy (anonymous):

then what

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

x^2+2x+k=x^2+10p+5x

OpenStudy (phi):

FOIL means first, outer, inner last (x+5)(x+p) First: x*x or x^2 outer: x*p or px inner: 5x last: 5p x^2 +px+5x+5p factor out the x to get x^2 +(p+5)x +5p set this equal to the other quadratic x^2 +(p+5)x +5p = x^2+2x+k now match coefficients: the "numbers" in front of the x must be equal. p+5 = 2 and 5p must equal k

OpenStudy (anonymous):

then k=-2x^2-7x-xp-5p

OpenStudy (phi):

it might not be obvious but when you have x^2 +(p+5)x +5p = x^2+2x+k it must be that x^2= x^2 (p+5)x = 2x 5p= k for this equation to always be true

OpenStudy (phi):

can you find p?

OpenStudy (anonymous):

yes

OpenStudy (phi):

once you find p, you can find k: 5p= k